OpenStudy (anonymous):

the larger of two numbers is five less than twice the smaller number. the sum of the two numbers is 16 find both numbers

OpenStudy (wolf1728):

A) L = (2 * S) -5 B) L + S = 16 A) (2*S) -5 -L =0 B) L + S =16 adding both equations 3S -5 = 15 3 S = 21 S=7 L = 16-S L = 9

OpenStudy (anonymous):

he/she was using algebra, where you substitute numbers for alphabets so you don't get confused when you solve equations i mean, it would be a little silly to put it as A) larger number = (2 * S) -5 B) larger number + smaller number = 16 A) (2*smaller number) -5 - larger number =0 B) larger number + smaller number =16 adding both equations 3(smaller number) -5 = 15 3(smaller number) = 21 and so on